# A Computational Theory for Efficient Model Evaluation with Causal Guarantees

### v3

@@ Abstract
In order to reduce the cost of experimental evaluation for models, we introduce a computational theory of evaluation for prediction and decision models: build evaluation model to accelerate the evaluation procedures. We prove upper bounds of generalized error and generalized causal effect error of given evaluation models. We also prove efficiency, and consistency to estimated causal effect from deployed subject to evaluation metric by prediction. To learn evaluation models, we propose a meta-learner to handle heterogeneous evaluation subjects space problem. Comparing with existed evaluation approaches, our (conditional) evaluation model reduced 24.1%-99.0% evaluation errors across 12 scenes, including individual medicine, scientific simulation, social experiment, business activity, and quantum trade. The evaluation time is reduced 3-7 order of magnitude comparing with experiments or simulations.
@@ 1 Introduction
Randomized controlled trials (Fisher (1974), Box et al. (2005)) are widely regarded as the gold standard for evaluating the efficacy of interventions, such as drugs and therapies, due to their ability to minimize bias and establish causal relationships between interventions and outcomes. By randomly assigning participants to intervention or control groups, RCTs effectively balance both observed and unobserved confounding variables, thereby mitigating the influence of hidden common causes (Reichenbach (1999)). However, conducting RCTs to assess every potential intervention is often prohibitively expensive and, in certain cases, unfeasible. This challenge is particularly pronounced in fields like artificial intelligence, where models may possess hundreds of parameters, resulting in an infinite and high-dimensional space of potential interventions. Evaluating such models through randomized experiments is resource-intensive, demanding substantial investments of time, finances, personnel, and materials (Speich et al. (2018)). Moreover, the extended duration of these experiments hampers the rapid iteration and optimization of models, rendering the process inefficient. Given these limitations, alternative methodologies that balance rigorous evaluation with practical feasibility are essential, especially in rapidly evolving domains requiring swift assessment and deployment of new models.
@@
An improvement is to build the centralized A/B platform, which has been proven successes in application updates and recommendation in Byte Dance (Byte-Dance ). The implementation of centralized A/B testing platforms can significantly reduce the costs associated with conducting randomized trials for researchers. However, this cost reduction does not inherently enhance the utility derived from each individual trial and may inadvertently increase the risk to participants. Notably, the lower cost and increased accessibility of such platforms can lead to a substantial rise in the number of randomized trials conducted. This proliferation of experiments raises ethical concerns regarding participant exposure to potential harm during the randomization process. For instance, as highlighted by Zhou et al. (2024b), the automation and scalability of A/B testing frameworks, while beneficial for rapid experimentation, necessitate careful consideration of ethical implications to safeguard participant well-being. Polonioli et al. (2023) emphasizes the need for ethical and responsible experimentation to protect users and society.
@@
Evaluatology (Zhan et al. (2024)) emerges as an independent filed in recent years. However, there is still a lack of a theory to help us find an effective and efficient evaluation system that satisfies the concerns and interests of the stakeholders. Here, we introduce a computational theory to evaluate the effect of given evaluation subjects. The main benefit to evaluate by computation is its efficiency and low cost. If we learned the connection between the experiment result and pre-experiment subjects, the evaluation cost will be reduced dramatically comparing with experiments.
@@
Figure 1: Procedure of computational evaluation. C is evaluation condition, S is evaluation subject, M is evaluation metric.
@@
The procedure of our computational evaluation is listed in figure 6. First, data of evaluation condition, evaluation subject, and evaluation metric are collected. If causal relationship is needed, the subject is required to be independently, randomly, identically sampled (IRIS) so that the causal effect from the subject to the metric can be bounded in the following steps. Next, evaluation model is learned from the collected data, and generalized evaluation error is upper bounded under independently, identically distributed error (IIDE) assumption. Also, the evaluation model is assumed conditional unbiased to bound the causal effect from subject to metric. Then, learned evaluation model is applied to the downstream applications once the upper bounds are low enough to satisfy the real need. The mentioned assumptions are assessed by testing their necessary conditions.
@@
In addressing on on the heterogeneous subject space problem in evaluation model learning, we propose a meta algorithm to parameterize subjects, learn (conditional) evaluation models, and inference separately. The upper bounds of generalized error for both metric and its first order difference on subject are minimized by minimizing the empirical error. Also, learning evaluation model from evaluation units directly is very challenging in some cases. So we add a proxy module for the subject in addressing the challenge. The proxy module will create some proxy metrics according some existing computational evaluation methods, which can help us to improve the performance of evaluation model.
@@
In our evaluation experiment, we test evaluation model on 11 scenes including individual medicine, scientific simulation, insurance, advertisement, trade. Comparing with the existed computational evaluation methods, our method reduced 24.1%-99.0% evaluation errors while remain the advantage of computational evaluation efficiency. We also test our conditional evaluation model in the quantitative trade scene. Comparing with baseline backtesting strategy, our evaluation model reduced 89.4% evaluation errors. The evaluation acceleration ratio is ranged from 1000 times to 10000000 times comparing with experimental or simulation based evaluation in the 12 scenes.
@@ 2 Notation
want to deploy, remove, understand, and modify
@@
Table 1 shows the notation we used in this paper. In order to simplify the evaluation problem, evaluation metric and difference of evaluation metric are assumed as measurable and pairwise comparable. The computational evaluation problem is modeled as computing the evaluation metrics for evaluation subjects with certain evaluation conditions.
@@ 3.1 Upper bound of generalized evaluation error
In Theorem 1, we show that the generalization error $E_{gen}$ is bounded by the empirical error $E_{emp}$. the number of error measurements $n$, and the significance level $1-\sigma$ under the IIDE assumption. This result underpins our ability to use a finite number of evaluation samples to compare the performance of evaluation models across an infinite subject and condition space. For example, when the evaluation subject space is extremely large or infinite—as in the cases of AI diagnosis models, AI treatment models, or quantitative strategy models—Theorem 1 allows us to generalize our empirical analysis to the entire subject space. Similarly, in evaluation condition spaces, such as those involving AI-driven discovery of unseen materials based on molecular and structural conditions, the generalization errors for any evaluation model can be bounded using the same theorem.
Once the upper bound is smaller than given error tolerance, evaluation model $f$ is regarded to meet the practical needs with high probability $1-\sigma$. Additionally, Theorem 1 provides a means to estimate the number of error measurements (or evaluation samples) required prior to sampling.
We emphasize that IIDE can also be satisfied if the non-IID components of $(C,S,M)$ was offset by the evaluation model or the evaluation error function although independent identical distributed $(C,S,M)$ is a sufficient condition of IIDE. In fact, the assumption of independent and identically distributed samples is considered excessively strong in this paper, particularly with respect to IID conditions and IID outcomes.
@@ 3.2 Hidden common cause
The hidden common cause between subject and metric may introduce the confounding bias if the evaluation task is causality sensitive, such as individual medicine. Causal inference is often used to handle hidden common cause from data. However, applying existed methods to handle hidden common cause problem in evaluation faced those challenges:
  • Hidden identical common cause (HIC). If the assignment of $S$ is determined by an unlisted identical common cause $S:=E$, and there exists effect from $E$ to $EI$. We may mistakenly attributing the effect on $EI$ of $E$ to $S$. The identical hidden confounder can never be excluded from statical analysis due to the identical property if $S$ was never controlled by randomization. Therefore, effect from $S$ to $EI$ is unidentifiable in all kinds of non-intervention based analysis, including diagram-based identification (Pearl (1995), Tian and Pearl (2002), Shpitser and Pearl (2006), Yan (2025)), observation study (ROSENBAUM and RUBIN (1983)), causal representation learning (Schölkopf et al. (2021)), and causal discovery (Spirtes and Glymour (1991), Zhang and Hyvärinen (2009)).
  • High-dimensional evaluation subject. Potential outcomes (Splawa-Neyman et al. (1990), Holland (1986), Imbens and Rubin (2015)) often assumes each evaluation subject must be sampled once which often do not established for high-dimensional evaluation subjects. For instance, parameters of AI model can be hundreds or thousands and more complex which structure is different from continuous dose-level analysis with low dimension.
For causality sensitive task, any hidden common cause between subject and metric is excluded including HIC if the subject was randomly, independently, identically sampled or assigned as shown in theorem 2, 3, and 4. For high-dimensional evaluation subject, the causal effect of evaluation error is bounded and evaluation result can be generalized to the whole subject space whatever how large the space size as shown in theorem 4.
@@
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled from distribution $Pr(S)$, $m$ is generate by a unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is root mean square error, $\forall f_{1}$, $\forall f_{2}$ where $f_{1}$ and $f_{2}$ are unbiased estimation of $f$ given $c$, we have Efficiency: if $E_{gen}(f_{1})<E_{gen}(f_{2})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(f_{1}(c,s_{a})-f_{1}(c,s_{b}))<E_{gen}(f_{2}(c,s_{a})-f_{2}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
From theorem 2, minimization of generalized mean square causal effect evaluation error can be performed by minimizing generalized evaluation error greedily until the evaluation error is close to 0.
@@
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, identically distributed error (IIDE) measurements where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, $s$ is independently, randomly, identically sampled (IRIS), evaluation error function $L$ is mean square error ranged from 0 to 1.
@@
Theorem 3 provides the rationale for why our evaluation models can learn the causal effect from subjects to metrics rather than mere correlations when the evaluation subjects are independently, randomly, and identically sampled (IRIS), the evaluation models are unbiased, and the evaluation errors for are independently, randomly, identically distributed (IIDE). In practice, $n_{a}$ and $n_{b}$ are typically zero for infinite evaluation subject (non-positivity Imbens and Rubin (2015)); in such cases, upper bound on the causal effect error is given by Theorem 4. Proofs of all the theorems and specific upper bounds were given in appendix A, G.
Comparing with Saito and Yasui (2020) and Alaa and Van Der Schaar (2019), our unbiased assumption has potential to be rejected from real data testing as shown in the experiment part while they did not provide approaches to reject their unbiased causal effect assumption (Saito and Yasui (2020)) or min-max optimal effect estimator assumption (Alaa and Van Der Schaar (2019)) from real data without counterfactual. Also, IID samples assumption in their work is too strong for causal evaluation, especially IID conditions and IID outcomes.
@@ 4 Learning evaluation models
In evaluation processes, two critical factors are evaluation error and evaluation cost. Our objective is to reduce evaluation cost while maintaining an acceptable level of generalized evaluation (causal) error through computational evaluation. To achieve this, we can minimize empirical evaluation error because upper bound size is fixed given number of error measurements $n$ and confidence requirement $1-\sigma$ in theorem 1.
Directly learning an evaluation model from evaluation samples presents significant challenges. To address this, we introduce a proxy module that generates surrogate metrics using established computational methods, such as dataset-based approaches including bootstrapping, hold-out validation, and cross-validation. These techniques enhance the performance of the evaluation model by providing more robust and reliable features. Furthermore, to effectively manage heterogeneous subjects, we employ vectorization strategies, develop specialized sub-models, and conduct inferences tailored to each subject type. This approach ensures that the evaluation model can accommodate the diverse characteristics inherent in heterogeneous subject spaces.
@@
Figure 2: Learn evaluation models from data.

### v4

# A Computational Theory for Efficient Mini Agent Evaluation with Causal Guarantees

@@ Abstract
In order to reduce the cost of experimental evaluation for agents, we introduce a computational theory of evaluation for mini agents: build evaluation model to accelerate the evaluation procedures. We prove upper bounds of generalized error and generalized causal effect error of given evaluation models for infinite agents. We also prove efficiency, and consistency to estimated causal effect from deployed agents to evaluation metric by prediction. To learn evaluation models, we propose a meta-learner to handle heterogeneous agents space problem. Comparing with existed evaluation approaches, our (conditional) evaluation model reduced 24.1% to 99.0% evaluation errors across 12 scenes, including individual medicine, scientific simulation, social experiment, business activity, and quantum trade. The evaluation time is reduced 3 to 7 order of magnitude per subject comparing with experiments or simulations.
@@ 1 Introduction
Randomized controlled trials (Fisher (1974), Box et al. (2005)) are widely regarded as the gold standard for evaluating the efficacy of interventions, such as drugs and therapies, due to their ability to minimize bias and establish causal relationships between interventions and outcomes. By randomly assigning participants to intervention group or control group, RCTs effectively balance both observed and unobserved confounding variables, thereby mitigating the influence of hidden common causes (Reichenbach (1999)). However, conducting RCTs to assess every potential agent is often prohibitively expensive and, in certain cases, unfeasible as shown in appendix A. This challenge is particularly pronounced in fields like artificial intelligences, where models may possess hundreds of parameters, resulting in an infinite and high-dimensional space of potential interventions. Evaluating such agents through randomized experiments is resource-intensive, demanding substantial investments of time, finances, personnel, and materials (Speich et al. (2018)). Moreover, the extended duration of these experiments hampers the rapid iteration and optimization of agents, rendering the process inefficient. Given these limitations, alternative methodologies that balance rigorous evaluation with practical feasibility are essential, especially in rapidly evolving domains requiring swift assessment and deployment of new agents.
@@
An improvement is to build the centralized A/B platform, which has been proven successes in application updates and recommendation in Byte Dance (Byte-Dance ). The implementation of centralized A/B testing platforms can significantly reduce the costs associated with conducting randomized trials for researchers. However, this cost reduction does not inherently enhance the utility derived from each individual trial and it still can not handle the infinite evaluation subjects problem. Notably, the lower cost and increased accessibility of such platforms can lead to a substantial rise in the number of randomized trials conducted. This proliferation of experiments raises ethical concerns regarding participant exposure to potential harm during the randomization process. For instance, as highlighted by Zhou et al. (2024b), the automation and scalability of A/B testing frameworks, while beneficial for rapid experimentation, necessitate careful consideration of ethical implications to safeguard participant well-being. Polonioli et al. (2023) emphasizes the need for ethical and responsible experimentation to protect users and society.
@@
In addition to other agent evaluation works in appendix B, evaluatology (Zhan et al. (2024)) emerges as an independent filed in recent years. However, there is still a lack of a theory to help us find an effective and efficient evaluation system that satisfies the concerns and interests of the stakeholders. Here, we introduce a computational theory to evaluate the effect of given agents. The main benefit to evaluate by computation is its efficiency and low cost. If we learned the connection between the experiment result and pre-experiment agents, the evaluation cost will be reduced dramatically comparing with experiments.
@@
Figure 1: Procedure of computational evaluation. C is evaluation condition, S is evaluation subject (agent), M is evaluation metric, EM is evaluation model.
@@
The procedure of our computational evaluation is listed in figure 7. First, data of evaluation condition, deployed agents, and evaluation metric are collected. If causal relationship is needed, the subject is required to be independently, randomly, identically sampled (IRIS) so that the causal effect from the subject to the metric can be bounded in the following steps. Next, evaluation model is learned from the collected data, and generalized evaluation error is upper bounded under independently, identically distributed error (IIDE) assumption. Also, the evaluation model is assumed unbiased to bound the causal effect from agent to metric. Then, learned evaluation model is applied to the downstream applications once the upper bounds are low enough to satisfy the real need. The mentioned assumptions are assessed by testing their necessary conditions.
@@
In addressing on on the heterogeneous agents space problem in evaluation model learning, we propose a meta algorithm to parameterize agents, learn (conditional) evaluation models, and inference separately. The upper bounds of generalized error for both metric and its first order difference on agents are minimized by minimizing the empirical error. Also, learning evaluation model from evaluation samples directly is very challenging in some cases. So we add a proxy module for the agents in addressing the challenge. The proxy module will create some proxy metrics according some existing computational evaluation methods, which can help us to improve the performance of evaluation model.
@@
In our evaluation experiment, we test evaluation model on 11 scenes including individual medicine, scientific simulation, insurance, advertisement, trade. Comparing with the existed computational evaluation methods, our method reduced 24.1% to 99.0% evaluation errors while remain the advantage of computational evaluation efficiency. We also test our conditional evaluation model in the quantitative trade scene. Comparing with baseline backtesting strategy, our evaluation model reduced 89.4% evaluation errors. The evaluation acceleration ratio is ranged from 1000 times to 10000000 times comparing with experimental or simulation based evaluation in the 12 scenes.
@@ 2 Notation
(agents) want to deploy, remove, understand, and modify
@@
Table 1 shows the notation we used in this paper. In order to simplify the evaluation problem, evaluation metric and difference of evaluation metric are assumed as measurable and pairwise comparable. The computational evaluation problem is modeled as computing the evaluation metrics for evaluation subject with certain evaluation conditions. Unless otherwise specified, all evaluation subjects in this paper are agent-based systems.
@@ 3.1 Upper bound of generalized evaluation error
In Theorem 1, we show that the generalization error $E_{gen}$ is bounded by the empirical error $E_{emp}$. the number of error measurements $n$, and the significance level $1-\sigma$ under the IIDE assumption. This result underpins our ability to use a finite number of evaluation samples to compare the performance of evaluation models across an infinite subject and condition space. For example, when the evaluation subject space is extremely large or infinite—as in the cases of AI diagnosis agents, AI treatment agents, or quantitative strategy agents—Theorem 1 allows us to generalize our empirical analysis to the entire subject space. Similarly, in evaluation condition spaces, such as those involving AI-driven discovery of unseen materials based on molecular and structural conditions, the generalization errors for any evaluation model can be bounded using the same theorem.
Once the upper bound is smaller than given error tolerance, evaluation model $f$ is regarded to meet the practical needs with high probability $1-\sigma$. Additionally, Theorem 1 provides a method to estimate the number of error measurements (or evaluation samples) required prior to sampling.
We emphasize that IIDE can also be satisfied if the non-IID components of $(C,S,M)$ was offset by the evaluation model or the evaluation error function (such as non-misspecified time-series model on time-series data) although independent identical distributed $(C,S,M)$ is a sufficient condition of IIDE. In fact, the assumption of independent and identically distributed samples is considered excessively strong in this paper, particularly with respect to IID conditions and IID outcomes.
@@ 3.2 Hidden common cause
The hidden common cause between subject and metric may introduce the confounding bias if the evaluation task is causality sensitive, such as individual medicine. Causal inference is often used to handle hidden common cause from data. However, applying existed methods to handle unmeasured hidden common cause problem in evaluation faced those challenges:
  • Hidden identical common cause (HIC). If the assignment of $S$ is determined by an unlisted identical common cause $E$ which can be denoted as $S:=E$, and there exists effect from $E$ to $M$. We may mistakenly attributing the effect on $M$ of $E$ to $S$. The identical hidden confounder can never be excluded from statistical analysis due to the identical property if $S$ was never controlled by randomization. Therefore, effect from $S$ to $M$ is unidentifiable in all kinds of existed non-intervention based causal analysis, including diagram-based identification (Pearl (1995), Tian and Pearl (2002), Shpitser and Pearl (2006), Bareinboim and Pearl (2012), YAN (2025)), neural-based identification (Xia et al. (2021)), statistical-based identification (Jaber et al. (2019)), observation study (ROSENBAUM and RUBIN (1983)), causal representation learning (Schölkopf et al. (2021)), and causal discovery (Spirtes and Glymour (1991), Zhang and Hyvärinen (2009)).
  • High-dimensional evaluation subject. Potential outcomes (Splawa-Neyman et al. (1990), Holland (1986), Imbens and Rubin (2015)) often assumes each evaluation subject must be sampled once which often do not established for high-dimensional evaluation subjects. For instance, parameters of AI model can be hundreds or thousands and more complex which structure is different from continuous dose-level analysis with very low dimension.
For causality sensitive task, any hidden common cause between subject and metric is excluded including HIC if the subject was randomly, independently, identically sampled or assigned as shown in theorem 2 and theorem 4. For high-dimensional evaluation subject, the causal effect of evaluation error is bounded and evaluation result can be generalized to the whole subject space whatever how large the space size as shown in theorem 4.
@@
Strict causal advantage. Given $\{c_{i},s_{i},m_{i}\}_{i=1}^{n}$, where $c$ is evaluation condition, $s$ is independent from $c$, and $e$, and $s$ is independently, randomly, identically sampled (IRIS) from distribution $Pr(S)$, $m$ is generate by an unknown and inaccessible function $m=f(c,e,s)$, $e$ is other unlisted random variable, evaluation error function $L$ is mean square error, $\forall\hat{f_{1}}$, $\forall\hat{f_{2}}$ where $\hat{f_{1}}$ and $\hat{f_{2}}$ are unbiased estimation of $f$ given $c$ (CU), we have Efficiency: if $E_{gen}(\hat{f_{1}})<E_{gen}(\hat{f_{2}})$ for any $s$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $E_{gen}(\hat{f_{1}}(c,s_{a})-\hat{f_{1}}(c,s_{b}))<E_{gen}(\hat{f_{2}}(c,s_{a})-\hat{f_{2}}(c,s_{b}))$; Consistency: if $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f})=0$, then $\forall s_{a}\in\mathbf{S}$, $\forall s_{b}\in\mathbf{S}$, $\forall c\in\mathbf{C}$, $\lim_{\hat{f}\rightarrow f}E_{gen}(\hat{f}(c,s_{a})-\hat{f}(c,s_{b}))=0$.
From theorem 2, minimization of generalized mean square causal effect evaluation error can be performed by minimizing generalized evaluation error greedily until the evaluation error is close to 0. However, we can not get the generalized evaluation errors in reality. An upper bound with empirical errors was inferred in the theorem 3.
@@
, where $\Delta\hat{f}=\hat{f}(c,s_{a})-\hat{f}(c,s_{b})$, $n_{a}$ and $n_{b}$ are number of independently, randomly, identically sampled error measurements (IIDE) where $s=s_{a}$ and $s=s_{b}$, $10<1-\sigma<$ 1 is confidence, evaluation error function $L$ is mean square error ranged from 0 to 1.
In practice, $n_{a}$ and $n_{b}$ are typically zero for infinite evaluation subject (non-positivity Imbens and Rubin (2015)); in such cases, upper bound of generalized effect error is given by Theorem 4.
@@
Theorem 4 provides the rationale for why our evaluation models can learn the causal effect from subjects to metrics rather than mere correlations when the evaluation subjects are independently, randomly, and identically sampled (IRIS), the evaluation models are unbiased, and the evaluation errors are independently, randomly, identically distributed (IIDE). Also, effect error is upper bounded by its bias even the model is not unbiased. Proofs of all the theorems were given in appendix C.
Comparing with Saito and Yasui (2020) and Alaa and Van Der Schaar (2019), our unbiased assumption has potential to be rejected from real data testing as shown in the experiment part while they did not provide approaches to reject their unbiased causal effect assumption (Saito and Yasui (2020)) or min-max optimal effect estimator assumption (Alaa and Van Der Schaar (2019)) from real data without counterfactual. Also, IID samples assumption in those works (Saito and Yasui (2020), Johansson et al. (2022), Alaa and Van Der Schaar (2019), Shalit et al. (2017), Li and Pearl (2022)) is too strong for causal evaluation, especially IID conditions and IID outcomes.
@@ 4 Learning evaluation models
In evaluation processes, two critical factors are evaluation error and evaluation cost. Our objective is to reduce evaluation cost while maintaining an acceptable level of generalized evaluation (causal) error through computational evaluation. To achieve this, we can minimize empirical evaluation error because upper bound size is fixed given number of error measurements $n$ and confidence requirement $1-\sigma$ in theorem 1 and theorem 4.
Directly learning an evaluation model from evaluation samples presents significant challenges. To address this, we introduce a proxy module that generates surrogate metrics using established computational methods, such as dataset-based approaches including bootstrapping, hold-out validation, and cross-validation. These techniques enhance the performance of the evaluation model by providing more robust and reliable features. Furthermore, to effectively manage heterogeneous agents, we employ vectorization strategies, develop specialized sub-models, and conduct inferences tailored to each agent type. This approach ensures that the evaluation model can accommodate the diverse characteristics inherent in heterogeneous agent spaces.
@@
Figure 2: Learn evaluation models from data. $P$ is proxy metrics of evaluation subject, and $T_{i}$ is tensorization function for subject type $i$.